# ATOM: A Pretrained Neural Operator for Multitask Molecular Dynamics

### v1

@@ 3.2.3 Training with Label Noise Regularization
Many DFT datasets are inherently noisy (Christensen & von Lilienfeld, 2020), and MD models can overfit to this noise. Motivated by the regularization effect of label noise (Damian et al., 2021; HaoChen et al., 2020), we augment the observed node positions $\mathbf{x}$ and velocities $\mathbf{v}$ by random Gaussian noise $\boldsymbol{\xi}_{\mathbf{x}},\boldsymbol{\xi}_{\mathbf{v}}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ during training. Let ${\mathcal{G}}^{(t)}_{\boldsymbol{\xi}}=(\mathbf{x}_{i}^{(t)}+\boldsymbol{\xi}_{x,i},\mathbf{v}^{(t)}+\boldsymbol{\xi}_{v,i})$ be the noised initial state at time $t$. We minimize the following regularized loss
@@ 3.3 Multitask ATOM Pretraining and TG80 Dataset
where we take expectation with respect to initial states of multiple molecules in the training set, as well as the time lags. Here, we emphasize ATOM also takes a time lag, $\Delta t$, as input.
TG80 Dataset. To facilitate pretraining of our neural operator, we introduce TG80, a superset of the MD17 dataset. The initial seed set comprises 40 molecules: 8 MD17 compounds and 32 additional drug-like molecules selected through expert review. We then augment the seed molecules with structurally similar molecules from the PubChem dataset of 173 million compounds (Bolton et al., 2011).
Accepted candidates had an ECFP-4 Tanimoto similarity between 0.875 and 0.925 to at least one seed molecule, and no more than 0.80 similarity to previously accepted molecules, alongside other criteria detailed in Section C.4 (Landrum et al., 2025; Rogers & Hahn, 2010; Rogers & Tanimoto, 1960). These thresholds follow common practice in the literature, balancing diversity while avoiding collapse into overly narrow chemical subspaces (Matter, 1997; Menke et al., 2021; Eastman et al., 2023; Harper et al., 2004; Zhang et al., 2023).
We generate all trajectories using ORCA V6.01 (Neese, 2022) with the PBE functional (Perdew et al., 1996), def2-SVP basis set (Weigend & Ahlrichs, 2005), $\Delta$4 dispersion corrections (Caldeweyher et al., 2019, 2020; Wittmann et al., 2024) at one femtosecond resolution, 300K temperature, in vacuum. This resembles an enhanced RMD17, with more modern dispersion corrections to improve stability and allow for a larger step size (Christensen & von Lilienfeld, 2020). As a result, TG80 exhibits more diverse dynamics and improved numerical stability, with no compound exceeding 50 Å center-of-mass drift in Figure 9111Simulations ran on 32 AMD EPYC 7543 cores with 256 GB RAM per molecule, totalling 806,400 CPU-hours (quoted market cost USD 150 000)..
@@ 4 Experiment Results
Baselines. For comparison, we include a range of classic to state-of-the-art baselines, including Radial Field (RF) (Köhler et al., 2019), Tensor Field Networks (TFN) (Thomas et al., 2018), SE(3) Transformer (SE(3)-Tr.) (Fuchs et al., 2020), E(n) equivariant graph neural networks (EGNN) (Garcia Satorras et al., 2021), and EGNO (Xu et al., 2024). Our EGNN baselines are EGNN-Rollout (EGNN-R), which predicts timesteps autoregressively, and EGNN-Sequential (EGNN-S), which uses the output of each GNN as the prediction of a given frame. We set all baseline hyperparameters following previous works (Xu et al., 2024, 2022; Shi et al., 2021) and tune ATOM and EGNO hyperparameters as in Table 16 and Table 17.
Training setups. For training of ATOM and EGNO, we consider two temporal discretization strategies in selecting the timestamps $t_{p}=t+\sum_{r=1}^{p}\Delta t_{r}$: (1) Uniform discretization selects $t_{p}=t+p/P\Delta T$ and (2) Tail discretization selects $t_{p}=t+\overline{\Delta}+p/P(\Delta T-\overline{\Delta})$ for a lag $\overline{\Delta}\in[0,\Delta T]$. In the main paper, we present experiment results with uniform discretization and include the results with tail discretization in Appendix E. We perform early stopping on the lowest S2S validation loss checkpoint and report results as mean $\pm 2\sigma$ over three training runs. All experiments are run on an NVIDIA® RTX 5080 with wall-clock time and FLOP utilization detailed in Table 14.
@@ 4.1 Single-task Learning
We benchmark on the MD17, RMD17, and MD22 DFT MD trajectory datasets (Chmiela et al., 2017; Christensen & von Lilienfeld, 2020; Chmiela et al., 2023). We partition the trajectories into train/validation/test splits of sizes 500/2000/2000, set $\Delta T=3000$ and $P=8$, and train for 2500 epochs following (Xu et al., 2024). For the performance on MD17 ( Table 1), we directly quote the results from (Xu et al., 2024) except for EGNO. We design ATOM to have six transformer blocks with a hidden size of 256.
MD17 and RMD17. As shown in Table 1, ATOM compares favorably with state-of-the-art (SOTA) baselines on MD17 dataset, yielding average reductions of 14.96% (S2S mean squared error (MSE)) and 8.3% (S2T MSE) on average [^2] We exclude benzene from the table due to the previously discussed high numerical noise.. In Table 9 (Appendix E.2), we benchmark ATOM on RMD17, and observe similarly competitive performance against EGNO.
@@
MD22. To evaluate performance on larger molecules, we consider Ac-Ala3-NHMe (20 heavy atoms), docosahexaenoic acid (DHA with 24 heavy atoms), and stachyose (45 heavy atoms) from the MD22 dataset (Chmiela et al., 2023). ATOM remains competitive on these systems; whereas EGNO fails to converge (Table 2). We attribute this discrepancy to differing inductive biases: GNNs such as EGNO restrict message passing to a predefined bond or radius graph and can therefore under-represent long-range, non-bonded steric and electrostatic interactions that dominate the behavior of large, sparsely connected molecules (Alon & Yahav, 2021; Kosmala et al., 2023). This explains the poor performance of EGNO on DHA, a prototypical sparse molecule (Nv et al., 2003) (Figure 3), compared to its relatively stronger results on the densely bonded stachyose. In contrast, the fully connected point-cloud attention in ATOM imposes no prior on interaction ranges, allowing simultaneous learning of both local and global interactions.
@@ 4.2 Multitask Learning on TG80
We pretrain ATOM on TG80, scaling to six attention blocks with a hidden size of 256. We select stochastic horizons $\Delta T\sim\operatorname{LogUnif}(8,24\,000)$ and use a five-fold, cluster-based cross-validation. Specifically, we compute ECFP-4 fingerprints (Rogers & Hahn, 2010), embed them using UMAP (McInnes et al., 2018), and apply agglomerative clustering (Ward, 1963) to partition compounds into ten disjoint clusters. The folds are then formed by holding out clusters, ensuring that the train/validation/test sets occupy distinct regions of chemical space. This cluster-wise protocol minimizes leakage and more closely reflects the prospective scientific setting in which models must generalize to unseen molecules. Cluster-based approaches present more challenging generalization problems than random splits or common chemical-scaffold-based splits (Guo et al., 2024). In Appendix E.3, we also consider pre-training on a standard random split of molecules.
Table 3 benchmarks ATOM by assessing both in-distribution (ID) and out-of-domain (OOD) S2T performance. For the in-distribution setting, we train, validate, and test on molecules from the same cluster. We observe that ATOM outperforms existing baselines by an average of 83.96% in terms of S2T MSE. We then assess out-of-domain (OOD) generalization performance by predicting the dynamics of unseen compounds drawn from disjoint clusters. Under OOD settings, ATOM nearly halves the S2T MSE of EGNO, with an average improvement of $39.74\%$ across five cluster splits. Notably, OOD ATOM beats ID EGNO performance in four of five folds. This striking zero-shot generalization, realized without any exposure to the test molecules, confirms that ATOM uniquely learns robust, transferable knowledge of molecular dynamics. In Appendix E.4, we show similar outperformance in S2S prediction. In Appendix F.2, we show that the significantly improved multitask performance comes with a modest overhead in training time and in inference latency.
@@
Gap $96.75\% +$$91.70\% +$$41.22\%$
Table 3: Multi-task S2T MSE $(\times 10^{-2})$ on TG80 across five UMAP cluster assignments.
@@
Gap $78.04\%$ $\% 88.0988.09\%$ $88.78\%$ $83.99\%$
OOD EGNO $45.95{\scriptstyle\pm 0.80}$ $115.43{\scriptstyle\pm 13.23}$ $151.74{\scriptstyle\pm 0.57}$ $63.90{\scriptstyle\pm 0.69} 11$$13.68{\scriptstyle\pm 2.50}$
@@ 4.3 Temporal Gap and Timestep Invariance Properties
$\Delta T$ Invariance. We evaluate the performance of ATOM with varying $\Delta T$. We compare ATOM, EGNO, and EGNN on S2T MSE by fixing $P=8$ and sweeping $\Delta T$ logarithmically from 10 to 10 000 on a pretrained in-distribution multitask model (on Cluster 1). In Figure 5, we show that ATOM maintains its extrapolation advantage across the range compared to EGNO, particularly at larger $\Delta T$. Ablating T-RoPE (NoPE) removes such advantage by exihibiting EGNO-like error trend with substantially higher MSE. This underscores T-RoPE’s role in stable time-gap extrapolation.
$P$ invariance. Following the discretization invariance in neural operators, we expect ATOM and EGNO models to show consistent MSE as $P$ varies under uniform discretization (Kovachki et al., 2021). Figure 5 confirms such a conjecture by showing that multitask ATOM pretrained at $P=8$ maintain constant S2T MSE as $P$ ranges from 4 to 24.
Figure 4: Pretrained (ID) multitask S2T MSE across $\Delta T$ values.
Figure 5: ATOM and EGNO are discretization invariant, showing stable S2T MSE.
@@ 4.4 Ablation studies
Table 4: S2T MSE ($\times 10^{-2}$) of a fixed input frame rotated and unrotated by an $SO(3)$ matrix.
@@
Gap $66.28{\scriptstyle\pm 26.32}$ $627.53{\scriptstyle\pm 926.53}$
Equivariant lifting. We assess the quasi-equivariant design against fully equivariant and non-equivariant alternatives. As shown in Figure 7, replacing the equivariant lifting introduced in Section 3.2.1 with standard linear layers (no equivariant lifting) markedly degrades the performance of ATOM, increasing S2T MSE by 22.48. Consistently, Section 4.4 shows that non-equivariant lifting is vulnerable under $SO(3)$ rotations of the trajectories, whereas the quasi-equivariant ATOM remains comparatively robust. Notably, the fully-equivariant variant of ATOM, described in appendix Section D.1, also underperforms ATOM in both single-task (Figure 7) and multitask (Figure 7) settings, with the gap exaggerated in the multitask setting. This aligns with recent findings in relaxed equivariance, suggesting that strict equivariance can restrict the model capacity and complicate the optimization process (Elhag et al., 2025).
Heterogeneous attention. We evaluate the contribution from heterogeneous temporal attention by substituting with a standard self-attention on the phase space features. This replacement reduces accuracy, increasing S2S MSE by 0.47.
@@ C.3 Numerical Stability
(c) TG80 dataset exhibits the lowest centre-of-mass drift among the evaluated MD datasets.
@@ C.4 TG80 Generation Algorithm
which is identical to the definition of the Jaccard similarity in this case (Rogers & Tanimoto, 1960).

### v2

@@ 3.2.3 Training with Label Noise Regularization
Many DFT datasets are inherently noisy (Christensen and von Lilienfeld, 2020), and MD models can overfit to this noise. Motivated by the regularization effect of label noise (Damian et al., 2021; HaoChen et al., 2020), we augment the observed node positions $\mathbf{x}$ and velocities $\mathbf{v}$ by random Gaussian noise $\boldsymbol{\xi}_{\mathbf{x}},\boldsymbol{\xi}_{\mathbf{v}}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ during training. Let ${\mathcal{G}}^{(t)}_{\boldsymbol{\xi}}=(\mathbf{x}_{i}^{(t)}+\boldsymbol{\xi}_{x,i},\mathbf{v}^{(t)}+\boldsymbol{\xi}_{v,i})$ be the noised initial state at time $t$. We minimize the following regularized loss
@@
Comparison to EGNO. We highlight that ATOM adopts different design choices compared to EGNO. First, EGNO is an EGNN operating on fixed bond connectivity, whereas ATOM uses an E(3)-equivariant lifting layer followed by globally connected point-cloud attention, which better handles long-range and sparsely bonded interactions. Second, EGNO is strictly equivariant end-to-end, while ATOM is quasi-equivariant, enforcing equivariance only in the lifting stage and relaxing it in deeper transformer layers, which our ablations show improves accuracy. Third, EGNO models time via Fourier temporal convolution, whereas ATOM uses Temporal RoPE, allowing translation-invariant handling of irregular time gaps and stronger temporal extrapolation. T-RoPE also uniquely allows modifying the time-horizon $\Delta T$ at inference by modulating the rotary phases (Section 3.2.2). Consequently, a pretrained ATOM can be evaluated at arbitrary $\Delta T$ values without retraining.
@@ 3.3 Multitask ATOM Pretraining and TG80 Dataset
where we take expectation with respect to initial states of multiple molecules in the training set, as well as the time lags. Here, ATOM also takes a time lag, $\Delta t$, as input, to modulate T-RoPE phase.
TG80 Dataset. To facilitate pretraining of our neural operator, we introduce TG80, a superset of the MD17 dataset. The initial seed set comprises 40 molecules: 8 MD17 compounds and 32 additional drug-like molecules selected through expert review. We then augment the seed molecules with structurally similar molecules from the PubChem dataset of 173 million compounds (Bolton et al., 2011). Accepted candidates had an ECFP-4 Tanimoto similarity between 0.875 and 0.925 to at least one seed molecule, and no more than 0.80 similarity to previously accepted molecules, alongside other criteria detailed in Section C.4 (Landrum et al., 2025; Rogers and Hahn, 2010; Rogers and Tanimoto, 1960). These thresholds follow common practice in the literature, balancing diversity while avoiding collapse into overly narrow chemical subspaces (Matter, 1997; Menke et al., 2021; Eastman et al., 2023; Harper et al., 2004; Zhang et al., 2023).
We generate all trajectories using ORCA V6.01 (Neese, 2022) with the PBE functional (Perdew et al., 1996), def2-SVP basis set (Weigend and Ahlrichs, 2005), $\Delta$4 dispersion corrections (Caldeweyher et al., 2019; 2020; Wittmann et al., 2024) at one femtosecond resolution, 300K temperature, in vacuum. This resembles an enhanced RMD17, with more modern dispersion corrections to improve stability and allow for a larger step size (Christensen and von Lilienfeld, 2020). As a result, TG80 exhibits more diverse dynamics and improved numerical stability, with no compound exceeding 50 Å center-of-mass drift in Figure 8111Simulations ran on 32 AMD EPYC 7543 cores with 256 GB RAM per molecule, totalling 806,400 CPU-hours (quoted market cost USD 150 000)..
@@ 4 Experiment Results
Baselines. For comparison, we include a range of classic to state-of-the-art baselines, including Radial Field (RF) (Köhler et al., 2019), Tensor Field Networks (TFN) (Thomas et al., 2018), SE(3) Transformer (SE(3)-Tr.) (Fuchs et al., 2020), E(n) equivariant graph neural networks (EGNN) (Garcia Satorras et al., 2021), MACE (Batatia et al., 2023), and EGNO (Xu et al., 2024). We note that MACE is pretrained on the authors’ own dataset, so this is not a strictly like-for-like comparison. Our EGNN baselines are EGNN-Rollout (EGNN-R), which predicts timesteps autoregressively, and EGNN-Sequential (EGNN-S), which uses the output of each GNN as the prediction of a given frame. We set all baseline hyperparameters following previous works (Xu et al., 2024; 2022; Shi et al., 2021) and tune ATOM and EGNO hyperparameters as in Table 19 and Table 20.
Training setups. For training of ATOM and EGNO, we consider two temporal discretization strategies in selecting the timestamps $t_{p}=t+\sum_{r=1}^{p}\Delta t_{r}$: (1) Uniform discretization selects $t_{p}=t+p/P\Delta T$ and (2) Tail discretization selects $t_{p}=t+\overline{\Delta}+p/P(\Delta T-\overline{\Delta})$ for a lag $\overline{\Delta}\in[0,\Delta T]$. In the main paper, we present experiment results with uniform discretization and include the results with tail discretization in Appendix E. We perform early stopping on the lowest S2S validation loss checkpoint and report results as mean $\pm 2\sigma$ over three training runs. All experiments are run on an NVIDIA® RTX 5080 with wall-clock time and FLOP utilization detailed in Section F.2.
@@ 4.1 Single-task Learning
We benchmark on the MD17, RMD17, and MD22 DFT MD trajectory datasets (Chmiela et al., 2017; Christensen and von Lilienfeld, 2020; Chmiela et al., 2023). We partition the trajectories into train/validation/test splits of sizes 500/2000/2000, set $\Delta T=3000$ fs and $P=8$, and train for 2500 epochs following (Xu et al., 2024). For the performance on MD17 ( Table 1), we directly quote the results from (Xu et al., 2024) except for EGNO. We design ATOM to have six transformer blocks with a hidden size of 256.
MD17 and RMD17. As shown in Table 1, ATOM compares favorably with state-of-the-art (SOTA) baselines on MD17 dataset, yielding average reductions of 14.96% (S2S mean squared error (MSE)) and 8.3% (S2T MSE) on average [^2] We exclude benzene from the table due to the previously discussed high numerical noise.. In LABEL:tab:rmd17_results (Appendix 8), we benchmark ATOM on RMD17, and observe similarly competitive performance against EGNO.
@@
MD22. To evaluate performance on larger molecules, we consider Ac-Ala3-NHMe (20 heavy atoms), docosahexaenoic acid (DHA with 24 heavy atoms), and stachyose (45 heavy atoms) from the MD22 dataset (Chmiela et al., 2023). ATOM remains competitive on these systems; whereas EGNO fails to converge (Table 2). We attribute this discrepancy to differing inductive biases: GNNs such as EGNO restrict message passing to a predefined bond or radius graph and can therefore under-represent long-range, non-bonded steric and electrostatic interactions that dominate the behavior of large, sparsely connected molecules (Alon and Yahav, 2021; Kosmala et al., 2023). This explains the poor performance of EGNO on MD22, which contains prototypically sparse molecules such as DHA, shown in Figure 2 (Nv et al., 2003). We further disentangle the role of connectivity from the use of attention by training a variant, ATOM-GATv2, in which our heterogeneous temporal attention is replaced by graph attention network v2 (GATv2) layers (Brody et al., 2022) operating on the same bond/radius graph as EGNO. ATOM-GATv2 still substantially underperforms the full ATOM model, indicating that the performance gains stem from the fully connected point-cloud interaction pattern rather than from attention alone.
@@
MACE $6.95\pm\scriptstyle{0.00}$ $2.06\pm\scriptstyle{0.00}$ $17.99\pm\scriptstyle{0.26}$ $0.72\pm\scriptstyle{0.00}$ $1.05\pm\scriptstyle{0.00}$ $6.44\pm\scriptstyle{0.00}$ $0.75\pm\scriptstyle{0.00}$
@@
MACE $5.06\pm\scriptstyle{0.00}$ $2.84\pm\scriptstyle{0.00}$ $16.09\pm\scriptstyle{0.03}$ $7\pm\scriptstyle{0.00} 0.55±0$$5\pm\scriptstyle{0.00}$ $3.26\pm\scriptstyle{0.00}$ $1.08\pm\scriptstyle{0.00}$
@@ 4.2 Multitask Learning on TG80
We pretrain ATOM on TG80, scaling to six attention blocks with a hidden size of 256. We select stochastic horizons $\Delta T\sim\operatorname{LogUnif}(8\text{ fs},24\,000\text{ fs})$ and use a five-fold, cluster-based cross-validation. Specifically, we compute ECFP-4 fingerprints (Rogers and Hahn, 2010), embed them using UMAP (McInnes et al., 2018), and apply agglomerative clustering (Ward, 1963) to partition compounds into ten disjoint clusters. The folds are then formed by holding out clusters, ensuring that the train/validation/test sets occupy distinct regions of chemical space. This cluster-wise protocol minimizes leakage and more closely reflects the prospective scientific setting in which models must generalize to unseen molecules. Cluster-based approaches present more challenging generalization problems than random splits or common chemical-scaffold-based splits (Guo et al., 2024). In Appendix E.2, we also consider pretraining on a standard random split of molecules.
Table 3 benchmarks ATOM by assessing both in-distribution (ID) and out-of-domain (OOD) S2T performance. For the in-distribution setting, we train, validate, and test on molecules from the same cluster. We observe that ATOM outperforms existing baselines by an average of 83.96% in terms of S2T MSE. We then assess out-of-domain (OOD) generalization performance by predicting the dynamics of unseen compounds drawn from disjoint clusters. Under OOD settings, ATOM nearly halves the S2T MSE of EGNO, with an average improvement of $39.74\%$ across five cluster splits. Notably, OOD ATOM beats ID EGNO performance in four of five folds. This striking zero-shot generalization, realized without any exposure to the test molecules, confirms that ATOM uniquely learns robust, transferable knowledge of molecular dynamics. In Appendix 11, we show similar outperformance in S2S prediction. In Appendix F.2, we show that the significantly improved multitask performance comes with a modest overhead in training time and in inference latency.
@@
[…]
Table 3: Multitask S2T MSE $(\times 10^{-2})$ on TG80 across five UMAP cluster assignments.
@@
[…]
EGNO $45.95{\scriptstyle\pm 0.80}$ $115.43{\scriptstyle\pm 13.23}$ $151.74{\scriptstyle\pm 0.57}$ $63.90{\scriptstyle\pm 0.69} 11$$13.68{\scriptstyle\pm 2.50}$
@@ 4.3 Temporal Gap and Timestep Invariance Properties
$\Delta T$ Invariance. We evaluate the performance of pretrained ATOM(at fixed $\Delta T=3000$) with varying $\Delta T$ at inference. We compare ATOM, EGNO, and EGNN on S2T MSE by fixing $P=8$ and sweeping $\Delta T$ logarithmically from 10 to 10 000 fs on an in-distribution (Cluster 1) multitask model. In Figure 4, we show that ATOM maintains its extrapolation advantage across the range compared to EGNO, particularly at larger $\Delta T$. Ablating T-RoPE (NoPE) removes this advantage by exhibiting an EGNO-like error trend with substantially higher MSE. This underscores T-RoPE’s role in stable time-gap extrapolation.
$P$ Invariance. Following the discretization invariance in neural operators, we expect ATOM and EGNO models to show consistent MSE as $P$ varies under uniform discretization (Kovachki et al., 2021). Figure 4 confirms such a conjecture by showing that multitask ATOM pretrained at $P=8$ maintain constant S2T MSE as $P$ ranges from 4 to 24 at inference.
Figure 3: Pretrained ($\Delta T=3000$, ID) multitask S2T MSE across varying $\Delta T$ values.
Figure 4: Pretrained ($P=8$) ATOM and EGNO are discretization invariant, showing stable S2T MSE.
@@ 4.4 Ablation studies
Table 4: S2T MSE ($\times 10^{-2}$) of a fixed input frame rotated and unrotated by an $\mathrm{SO}(3)$ matrix.
@@
Increase $66.28{\scriptstyle\pm 26.32}$ $627.53{\scriptstyle\pm 926.53}$
[…]
Equivariant lifting. We assess the quasi-equivariant design against a non-equivariant ATOM. As shown in Figure 6, replacing the equivariant lifting introduced in Section 3.2.1 with standard linear layers (no equivariant lifting) markedly degrades the performance of ATOM, increasing S2T MSE by 22.48. Section 4.4 further quantifies sensitivity to $\mathrm{SO}(3)$ rotations: ATOM’s S2T MSE increases by $10.80\times$ under rotation, compared to $19.77\times$ without equivariant lifting. Notably, the fully-equivariant variant of ATOM, described in appendix Section D.1, also underperforms ATOM in both single-task (Figure 6) and multitask (Figure 6) settings, with the gap exaggerated in the multitask setting. This aligns with recent findings on relaxed equivariance, suggesting that strict equivariance can limit model capacity and complicate the optimization process (Elhag et al., 2025). We present estimates of the quasi-equivariance $\varepsilon$ in Section E.3.
Heterogeneous attention. We find that substituting heterogeneous temporal attention with standard self-attention on the phase space features increases S2S MSE by 0.47, suggesting that cross-attention enables access to non-trivial feature interactions.
@@ C.3 Numerical Stability
(c) TG80 molecules exhibit the lowest centre-of-mass drift among the evaluated MD datasets.
@@ C.4 TG80 Generation Algorithm
which is identical to the definition of the Jaccard similarity in this case (Rogers and Tanimoto, 1960).